(a) According to a certain chemical theory, the percentage of iron in a certain compound should be 12.1. To test this theory, it was decided to analyze 9 different random samples of the compound to see if the measurements would differ significantly from 12.1%. Suppose the 9 samples gave the following results:

11.7 12.2 10.9 11.4 11.3 12.0 11.1 10.7 11.6

Use either classical or p-value approach to verify the hypothesis at level of significance a= 0.05. Assume that the population is normally distributed.


(b) A bank manager estimates that 75% of the customers wait less than ten minutes in the queue at the bank. To look into this estimation, a customer group conducted a survey with a sample of 90 customers. It was found that 15% of the customers waited at least ten minutes in the queue. At the level of significance a= 0.01, is there enough evidence to conclude that more than 75% of the customers wait less than ten minutes in the queue at the bank?


1
Expert's answer
2021-11-23T13:30:36-0500

a.

For this question, we shall use the classical approach.

We first determine the value of the sample mean and standard deviation.

The mean is given by,

"\bar{x}=\sum(x)/n" where "n=9"

"\bar{x}=102.9/9=11.43333333"

To find the standard deviation, we need to find the variance given by the formula,

"V(X)=(\sum(x^2)-(\sum(x))^2/n)/(n-1)"

"V(X)=(1178.45-1176.49)\/8=0.245"

The standard deviation "SD(X)=\sqrt{V(X)}=\sqrt{0.245}=0.494974747"

The hypotheses tested are,

"H_0:\mu=12.1\space Against\space H_1:\mu\not=12.1"

To perform this test, we shall apply the student's t distribution since the population variance is not known and the sample size is less than 30.

The test statistic is given as,

"t^*=(\bar{x}-\mu)/(SD(X)/\sqrt{n})"

"t^*=(12.1-11.4333)/(0.494974747/\sqrt{9})=0.666667/0.164992=4.04061"

"t^*" is compared with the table value at "\alpha=0.05" with "(n-1)=9-1=8" degrees of freedom.

Thus, "t_{\alpha/2,8}=t_{0.05/2,8}=t_{0.025,8}=2.306", and the null hypothesis is rejected if "t^*\gt t_{0.025,8}"

Since "t^*=4.04061\gt t_{0.025,8}=2.306", we reject the null hypothesis and we conclude that there is sufficient evidence to show that the measurements differ significantly from 12.1% at 5% significance level.


b.

For this question, we test hypothesis on a single population proportion.

We test the following hypotheses,

"H_0:p=0.75"

"Against"

"H_1:p\gt 0.75"

From the information above, the proportion estimate "(\hat{p})" is 85%=0.85. We shall apply the z-test to perform this hypothesis test since "\hat{p}\sim N(p,(p*(1-p))/n)" where "p=0.75,n=90"

Now, the test statistic is given as,

"Z=(\hat{p}-p)/\sqrt{p*(1-p)/n}=(0.85-0.75)/\sqrt{(0.75*0.25)/90}=0.1/0.04564=2.1989(4dp)"

"Z" is compared with the table value at "\alpha=0.01".

The table value is,

"Z_{\alpha}=Z_{0.01}=2.33" and the null hypothesis is rejected if "Z\gt Z_\alpha"

Since "Z=2.1989\lt Z_\alpha=2.33", we fail to reject the null hypothesis and conclude that there is no enough evidence to show that more than 75% of the customers wait less than ten minutes in the queue at the bank at 1% level of significance.